OpenStudy (anonymous):

consider the line whose equation is 4x-8y=8 A.) write the equation in slope-intercept form B.) write a slope of a line perpendicular to this line C.) write the equation of the line, in point-slope form, that is perpendicular to the line that passes through the point (1, 2)

jigglypuff314 (jigglypuff314):

slope-intercept form -> y = mx + b so you want to isolate y from 4x-8y=8

OpenStudy (anonymous):

A.) I have y=1/2x-1 B.) I have y=2/1-1

jigglypuff314 (jigglypuff314):

the first is correct mmm for the second if you meant (2/1)(-1) then yes :)

OpenStudy (anonymous):

is there still the y=

jigglypuff314 (jigglypuff314):

slope is represented by the letter m so m = -2